1 L flask contains air, water vapour and small amount of liquid water at a pressure of 200 mm Hg. If this is connected to another 1 L evacuated flask, what will be the final pressure of the gas mixture t equilibrium. Assume the temp to be 50°C.(Aqueous tension at 50°C=93mm hg)
 
  1. 120.56 mm
  2. 230 mm
  3. 146.5 mm
  4. 109.4mm
      

Grade:Select Grade

1 Answers

User
6 years ago
Here the initial pressure is p1=200mm+93mm=293mm
let the final pressure be p2
now... convert the temperature to kelvin scale,t1=t2=323
p1*v1/t1=p2*v2/t2
293*1/323=p2*2/323
solving we get
p2=293/2
p2=146.5mm